OpenStudy (fanduekisses):

How would you integrate 8cos (pix)?

OpenStudy (fanduekisses):

I know the antiderivative of cos (x) is sin (x)

OpenStudy (fanduekisses):

-sinx? What do you mean?

OpenStudy (datman21328):

The anti derivative of cosx is -sinx

OpenStudy (datman21328):

Sorry I'm using mobile

OpenStudy (fanduekisses):

No lol

Nnesha (nnesha):

\[\large\rm \int\limits_{ }^{} 8\cos(\pi x)~ dx \] like this ?

OpenStudy (fanduekisses):

You're thinking about the derivative

OpenStudy (fanduekisses):

Yes

Nnesha (nnesha):

find anti derivative and since there is a funcation inside a function `pi x` you have to apply the chain rule

OpenStudy (fanduekisses):

Ohh i think there is something called substitution rule

OpenStudy (amtran_bus):